OpenStudy (horsegirl27):

HELP NEEDED MEDALS WILL BE GIVEN PLEASE HELP!!! I WILL BECOME YOUR FAN AND GIVE A MEDAL PLEASE PLEASE HELP ASAP!!!!! question will be posted below

OpenStudy (horsegirl27):

Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: A.f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 B.f(n) = 35n I.Which one of these formulas show the sequence written recursively, and which shows it written explicitly? Justify your explanations. II.Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means. III.Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. IV. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function. I really need help with part III I have the rest done just need help with III I GIVE MEDALS!!!!!!!!!1

OpenStudy (horsegirl27):

III.Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations

OpenStudy (horsegirl27):

That's all I need help with.

OpenStudy (horsegirl27):

@Daniellelovee @LetsLearn2000 @ILOVEHORSES

OpenStudy (horsegirl27):

@SolomonZelman

OpenStudy (horsegirl27):

please help I will give a medal

OpenStudy (horsegirl27):

@Daniellelovee @LetsLearn2000 please help me

OpenStudy (horsegirl27):

@Daniellelovee please help me again your great

OpenStudy (horsegirl27):

please help...

OpenStudy (horsegirl27):

@lexie<3 help me

OpenStudy (daniellelovee):

im trying but im so lazy and this is difficult lol

OpenStudy (horsegirl27):

ok thanks so much. Doesn't help by saying it's difficult lol. And me too haha but I really do need help and soon please

OpenStudy (horsegirl27):

As long as you try and help me learn it even if u yourself doesn't figure it out. I just need an answer so I can submit my homework

OpenStudy (horsegirl27):

So your working on it? ok

OpenStudy (daniellelovee):

yes lol give me time

OpenStudy (horsegirl27):

ok thanks so much

OpenStudy (anonymous):

I would help but that's really confusing, im soory

OpenStudy (daniellelovee):

can you post your calculations for the problems before? because that would help

OpenStudy (horsegirl27):

medals will be given!!! ok

OpenStudy (horsegirl27):

A: f(2)=f(1)+f(1)=70 shows it written recursively and the other shows it explicitly. B: To make a table I used this rule: f(1) = 35, f(n) = f(1) + f(n-1) for n > 1. The rule says to use n>1 but instead of 1, n is 2. n | f(n) ----------- 1 | 35 2 | ? f(n) = f(1)+ f(n-1) f(2) = f(1)+f(1) f(2) = 35+35 f(2) = 70 C: D: This sequence is a function because no numbers repeat and it goes up a certain amount each time.

OpenStudy (horsegirl27):

I need C (III)

OpenStudy (daniellelovee):

oh it works the same way as for f(2) except that multiply 35x40 this time at least thats what im understanding

OpenStudy (horsegirl27):

ok can you show me

OpenStudy (horsegirl27):

I just want to be able to check my own work. I have like no confidence in algebra sometimes. So if I can see what you did that would be GREAT just seeing it helps me build confidence

OpenStudy (daniellelovee):

ok if f(1)=35 and f(2)=70 then he pattern is multiply the n of f f(n) times 35 so F(40)=35x40=

OpenStudy (horsegirl27):

so 1400?

OpenStudy (daniellelovee):

thats what im getting from it

OpenStudy (horsegirl27):

f(40)=35*40 and 35*40=1400 so f(40)=1400

OpenStudy (daniellelovee):

yup

OpenStudy (horsegirl27):

cool

OpenStudy (daniellelovee):

just explain how you got there which would be by saying explaning the pattern of multipliying by 35

OpenStudy (horsegirl27):

ok thanks I just needed to be guided through it. Or I sort of panic and freak out and can't do it lol. Then I pace around cuddling with my little fluffy puppy then I tell my mom it's done then I go ride then I come back and I'm like omg it's not done omgomgomg what do I do??? True story haha

OpenStudy (horsegirl27):

Thank you soo much!

OpenStudy (daniellelovee):

haha i know lol we all have a subject that is frustrating and infuriating so we procastinate on it

OpenStudy (daniellelovee):

no problem :)

OpenStudy (horsegirl27):

Yes. And that is how I'm over a month behind since September. Freaking out, then procrastinating, more freaking out... lol. But my mom says if I catch up soon I get to fly to Florida with some horse friends and go to some big horse shows and of course, since I will be at West Palm Beach, the beach!

OpenStudy (horsegirl27):

Flown once from NC to my new house in NY when I was 6 have been to 1 big show other than the big stuff my mom and sis did when I was younger and not showing, then I got a baby to train so we are only just starting to do little shows, and I have never been to Florida so omg a great inspiration!

OpenStudy (horsegirl27):

baby horse.

OpenStudy (horsegirl27):

A dapple gray Arabian mare by Showkayce! Her name is Show Mia Rose

OpenStudy (horsegirl27):

barn name mia and her mom was named rose. Her mom was my sister's old Arab pony...